OpenStudy (nickymail48):

20x^3+49x^2+11x factor

OpenStudy (solomonzelman):

can you (at first) take an x out? Tell me what you get after doing that.

OpenStudy (nickymail48):

20x+49 i dont get the factoring

OpenStudy (solomonzelman):

\(\large\color{black}{ 20x^3+49x^2+11x}\) each term is divisible by x, correct?

OpenStudy (solomonzelman):

\(\large\color{black}{ 20x^3\div x = ?}\) \(\large\color{black}{ 49x^2\div x = ?}\) \(\large\color{black}{ 11x\div x = ?}\)

OpenStudy (nickymail48):

yes

OpenStudy (solomonzelman):

can you tell me what do you get for each term, when you divide it (every term) by x?

OpenStudy (nickymail48):

the same number

OpenStudy (solomonzelman):

\(\large\color{black}{ 20x^3+49x^2+11x}\) would be, \(\large\color{black}{ x(20x^2+49x+11)}\) because when you divide \(\large\color{black}{ 20x^3+49x^2+11x}\) you get, \(\large\color{black}{ 20x^2+49x+11}\) and multiply that times x on the outside, not to change the value.

OpenStudy (solomonzelman):

Sorry, I suck at explaining these...

OpenStudy (nickymail48):

no they hard! its ok thanks

OpenStudy (solomonzelman):

you can think of a way to factor the, \(\large\color{black}{ 20x^2+49x+11}\)

OpenStudy (solomonzelman):

\(\large\color{black}{ (5x+~~~~)(4x+~~~~)}\) what 2 numbers give you a product of 11, and sum of 49?

OpenStudy (solomonzelman):

(they have to be factors of 11)

OpenStudy (nickymail48):

7, 11, nothng go into 11

OpenStudy (solomonzelman):

only 1 and 11.

OpenStudy (nickymail48):

ok

OpenStudy (solomonzelman):

\(\large\color{black}{ (5x+11)(4x+1)}\)

OpenStudy (solomonzelman):

because 4*11=44 and 5*1=5, and 44+5=49 There is the middle term, So, \(\large\color{black}{ x(20x^2+49x+11)}\) becomes, \(\large\color{black}{ x(5x+11)(4x+1)}\)

OpenStudy (nickymail48):

so why is there x on outside of problem?

OpenStudy (solomonzelman):

because it was, \(\large\color{black}{ 20x^3+49x^2+11x}\) we factored out of x, \(\large\color{black}{ x(20x^2+49x+11)}\) and then factored the parenthesis, to get, \(\large\color{black}{ x(4x+1)(5x+11)}\)

OpenStudy (nickymail48):

ohh okthanks

OpenStudy (solomonzelman):

Yw